OpenStudy (anonymous):

Solve the system by substitution. -4.5x - 2y = -12.5 3.25x - y = -0.75

OpenStudy (anonymous):

Solve one of the equations for either x or y and then plug that into the other equation.

OpenStudy (anonymous):

Which one do you want to solve and for either x or y? Hi ( hint look for something easy like x or y with no value in front of it)

OpenStudy (anonymous):

X seems the easiest. times both sides by 2 to take out all the Ys?

OpenStudy (phoenixfire):

Are you still needing help @sage234 ?

OpenStudy (anonymous):

Yes @PhoenixFire I am horrible at this and cannot help but second guess myself

OpenStudy (anonymous):

@mariahsilva15 Can you help?

OpenStudy (anonymous):

Sage234 Privat message me and i will so help you (:

OpenStudy (anonymous):

 Make a x and y table.. that is the first step

OpenStudy (phoenixfire):

Rearrange the second equation so you have y=...... Show me what you get.

OpenStudy (phoenixfire):

\(-4.5x - 2y = -12.5\) <- EQ (1) \(3.25x - y = -0.75\) <- EQ (2) If you take (2) and rearrange for y.\[y=3.25x+0.75\]Now substitute this into (1) and solve for x\[-4.5x - 2(3.25x+0.75)=-12.5\] Can you solve this equation?